# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ 1.3 Discussion
The forgetful functor ${\mathrm{rep}}(B)\rightarrow{\mathop{\sf{Coh}}\nolimits}(G/B))$ is at the core of Kuznetsov-Polishchuk’s construction of exceptional collections on maximal Grassmannians, [34]. Starting off from a full exceptional collection in the equivariant category and suitably mutating the collection, they arrive at an exceptional collection of objects ${\mathcal{E}}^{\lambda}$ in the non-equivariant derived category where $\lambda$ runs over a certain collection of weights in $X(T)$. The relation of their construction to the one of the present paper is a subject of further research. We note that the objects ${\mathcal{X}}_{p}$, $p\in W$ are amenable to calculations; thus, given a group $G$ and a maximal parabolic subgroup $P\subset G$, the first step in relating the objects ${\mathcal{X}}_{p}$ to the objects ${\mathcal{E}}^{\lambda}$ of [34] is to compute the direct images $\pi_{\ast}{\mathcal{X}}_{p}$, $p\in W$ along the projection $\pi:G/B\rightarrow G/P$.
@@
Furthermore, the objects ${\mathcal{E}}^{\lambda}$ of [34] are a priori complexes of coherent sheaves. Quoting from loc.cit.: “Note that a priori the exceptional objects ${\mathcal{E}}^{\lambda}$ constructed above are complexes. However, we have the following Conjecture 4.1: For any exceptional block $B\subset P_{L}^{+}$ and $\lambda\in B$ the object ${\mathcal{E}}^{\lambda}$ is a vector bundle.” At the moment, it is unclear whether the objects ${\mathcal{X}}_{p}$, $p\in W$ are pure with respect to the standard $t$-structure on $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(G/B))$. Evidence for their purity comes from the rank two cases, see Section 14. It is natural to conjecture that ${\mathcal{X}}_{p}$, $p\in W$ are always pure.
@@
A very interesting question is to understand a relation of our results to baric structures/staggered $t$-structures on the derived categories of equivariant coherent sheaves from [1] and [2]. Further, full exceptional collections on flag varieties appear through the computations related to the dual Steinberg basis of ${\sf K}^{0}_{G}(G/B)$, [14]. These, in turn, are related to Lusztig’s asymptotic affine Hecke algebra and to Lusztig’s canonical basis in the $T$-equivariant $\sf K$-theory of Springer fibres. Ultimately, the problem is to relate the main construction of the objects $X_{p}$, $p\in W$ of the present paper to Lusztig’s canonical basis. Finally, the full exceptional collections in rank two from Section 14 possess a property that their duals are obtained from the initial ones by applying the Grothendieck duality followed by tensoring with the square root of the canonical sheaf. In particular, the dual collections in rank two also consist of pure objects. This property allows to put a highest weight category structure on the exact category of vector bundles ${\sf Vect}(G/B)$ in rank two cases and seems to be a manifestation of Ringel duality. Thus, a natural question is to see whether that property holds in general case; in particular, this means computing the dual collection to ${\mathcal{X}}_{p}$, $p\in W$.
It could be more economical to state some of the principal results in the language of quotient stacks; e.g., the passage from the semiorthogonal decomposition of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ as a $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$-linear category to that of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(G/B))$ as a ${\sf k}$-linear category could be deduced from the base change theorem for semiorthogonal decompositions, [33]. In the interests of full transparency of the arguments and for convenience of the reader, we have resorted to the more classical language in formulating and proving our results.
Notation
1.4 Semisimple groups
@@
Let ${\mathbb{G}}\rightarrow\mathbb{Z}$ be a simply connected semisimple Chevalley group scheme (a smooth split affine group scheme over ${\rm Spec}(\mathbb{Z})$ whose geometric fibres are connected simply connected semisimple algebraic groups), and ${\mathbb{G}}/{\mathbb{B}}\rightarrow\mathbb{Z}$ be the corresponding Chevalley flag scheme (resp., the corresponding parabolic subgroup scheme ${\mathbb{G}}/{\mathbb{P}}\rightarrow{\mathbb{Z}}$ for a standard parabolic subgroup scheme ${\mathbb{P}}\subset\mathbb{G}$ over ${\mathbb{Z}}$). Then ${\mathbb{G}/\mathbb{P}}\rightarrow{\rm Spec}({\mathbb{Z}})$ is flat and any line bundle $\mathbb{L}$ on ${\mathbb{G}/\mathbb{P}}$ also comes from a line bundle on ${\mathbb{G}/\mathbb{B}}$.
@@
Given an affine algebraic group $H$, flat over a noetherian base ring ${\sf k}$, an $H$-module is a ${\sf k}[H]$-comodule (a rational representation). The abelian category of $H$-modules is denoted ${\mathrm{Rep}}(H)$ and ${\mathrm{rep}}(H)$ denotes the full abelian subcategory of representations which are finitely generated over the base ring ${\sf k}$. (cf. [3, Section 1]).
@@
Let $G$ be obtained from ${\mathbb{G}}$ by base change along ${\rm Spec}({\sf k})\rightarrow{\rm Spec}(\mathbb{Z})$. When ${\sf k}={\mathbb{Z}}$ this just means that $G=\mathbb{G}$.
@@
Now let ${\sf k}$ be a field of arbitrary characteristic. So $G$ is a split semisimple simply connected affine algebraic group. Let $T$ be a split maximal torus of $G$, obtained by base change from a split maximal torus $\mathbb{T}$ in $\mathbb{B}$. Let $W=N_{G}({T})/T$ be the Weyl group and let $B\supseteq T$ correspond with $\mathbb{B}\supseteq\mathbb{T}$. Let $X=X(T)$ be the weight lattice and $\Phi$ be the root system. We choose the system $\Phi^{+}$ of positive roots which makes $B$ the negative Borel and let $\Pi$ be the set of simple roots. Let (,)(~{},~{})( , ) be a non-singular, symmetric, positive definite $W$-invariant form on ${\mathbb{R}}\otimes_{\mathbb{Z}}X$ and let $X(T)_{+}=\{\lambda\in X:(\lambda,\alpha^{\vee})\geq 0$ for all $\alpha\in S$}, be the set of dominant weights, where $\alpha^{\vee}=\frac{2\alpha}{(\alpha,\alpha)}$ for $\alpha\in\Phi$.
@@
We want to emphasize once again the fact that our convention (which is that of [29]) is that ${B}$ corresponds to the negative roots (as opposed to the conventions of [25], [41], and [46]). For that reason, the translation of various results from loc.cit. that are used extensively in this paper requires some care.
@@
Let $\Pi=\{\alpha_{1},\dots,\alpha_{r}\}$ be the set of simple roots relative to our choice of $\Phi^{+}$. The fundamental weights $\omega_{i}$ satisfy $(\omega_{i},\alpha_{j}^{\vee})=\delta_{ij}$. A parabolic subgroup of $G$ is usually denoted by $P$; in particular, for a simple root $\alpha\in\Pi$, denote ${P}_{\alpha}$ the minimal parabolic subgroup of $G$ associated to $\alpha$.
Given a $B$-module $M$ we associate to it as in [29, I 5.8] a sheaf $\mathcal{L}(M)$ on $G/B$ with fibre $M$ at the point $B/B$. Given a weight $\lambda\in{\rm X}({T})$, denote $\mathcal{L}_{\lambda}$ the corresponding line bundle on ${G}/{B}$. Here we still follow [29], so that $\mathcal{L}_{\lambda}$ or $\mathcal{L}(\lambda)$ is associated to the one dimensional $B$-module ${\sf k}_{\lambda}$ of weight $\lambda$, not associated to the dual of that representation.
@@
The weight lattice $X$ has a natural partial order $\geq_{d}$, known as the dominance order: for $\lambda,\mu\in X$ we write $\lambda\geq_{d}\mu$ if $\lambda-\mu$ is a sum of positive roots, with repetitions allowed [^2] That is, $\lambda-\mu$ is of the form $\sum_{\alpha\in\Phi^{+}}n_{\alpha}\alpha$, for non-negative integers $n_{\alpha}$.. We denote $w_{0}$ the longest element of the Weyl group $W$ and let $\lambda^{\ast}=-w_{0}\lambda$ for $\lambda\in X(T)_{+}$ (dual or contragredient, cf. [28, Section 2.2]).
@@
For $\lambda\in X$ let $\nabla_{\lambda}$ or $\nabla(\lambda)$ be the induced module ${\mathrm{ind}}_{B}^{G}({{\sf k}}_{\lambda})$ [29, I, Section 3.3]. It is finite-dimensional and non-zero if and only if $\lambda$ is dominant. For $\lambda\in X(T)_{+}$ we denote by $\Delta(\lambda)$ the Weyl module $(\nabla_{\lambda^{\ast}})^{\ast}$. Then $\nabla_{\lambda}$ has simple socle $L(\lambda)$ and $\Delta(\lambda)$ has simple head $L(\lambda)$. If ${\sf k}={\mathbb{Z}}$ then $\nabla_{\lambda}$, $\Delta(\lambda)$ are finitely generated and free over ${\sf k}$, but heads and socles make less sense.
@@
([29, II Proposition 4.13, B.4]) Let $\lambda\in X(T)_{+}$.
${\mathrm{Ext}}^{i}_{G}({\sf k},\nabla(\lambda))={\mathrm{Ext}}^{i}_{B}({\sf k},{\sf k}_{\lambda})=\left\{\begin{array}[]{l}{\sf k},\quad{\rm for}\quad i=0,\ \lambda=0,\\ 0,\quad{\rm otherwise}.\\ \end{array}\right.$
1.5 Triangulated categories
@@
2 Rappels: $B$-modules
2.1 Highest weight category structure on ${\mathrm{Rep}}(G)$
@@
Recall that relative to the choice of the system of positive roots $\Phi^{+}$ there is a partial order $\leq_{d}$ on $X(T)$: given two weights $\lambda,\mu\in X(T)$ we write $\lambda\leq_{d}\mu$ if $\mu-\lambda$ is a nonnegative integer combination of elements of $\Phi^{+}$. Simple modules in ${\mathrm{Rep}}(G)$ are parametrized by the dominant weights $X(T)_{+}$, [29, II, Section 2]. By [13] and [17], there is a highest weight category structure on ${\mathrm{Rep}}(G)$ with the weight poset $(X(T)_{+},\leq_{d})$: the standard modules in this structure are the Weyl modules $\Delta_{\lambda},\lambda\in X(T)_{+}$ and the costandard modules are the dual Weyl modules $\nabla_{\lambda},\lambda\in X(T)_{+}$.

### v3

@@ 1.3 Discussion
The forgetful functor ${\mathrm{rep}}(\mathbf{B})\rightarrow{\mathop{\sf{Coh}}\nolimits}(\mathbf{G}/\mathbf{B})$ underlines Kuznetsov–Polishchuk’s construction of exceptional collections on maximal Grassmannians, [37]. Starting off from a full exceptional collection in the equivariant category and suitably mutating the collection, they arrive at an exceptional collection of objects ${\mathcal{E}}^{\lambda}$ in the non-equivariant derived category where $\lambda$ runs over a certain collection of weights in $X(\mathbf{T})$. It is an interesting question to understand the relation of the objects ${\mathcal{E}}^{\lambda}$ from loc.cit. to the the objects $\hat{\mathcal{X}}_{v},v\in W^{\mathbf{P}}$ from Theorem 13.5.
@@
Furthermore, the objects ${\mathcal{E}}^{\lambda}$ of [37] are a priori complexes of coherent sheaves. Quoting from loc.cit.: “Note that a priori the exceptional objects ${\mathcal{E}}^{\lambda}$ constructed above are complexes. However, we have the following Conjecture 4.1: For any exceptional block $B\subset P_{L}^{+}$ and $\lambda\in B$ the object ${\mathcal{E}}^{\lambda}$ is a vector bundle.” At the moment, it is not clear whether the objects ${\mathcal{X}}_{p}$, $p\in W$ from Theorem 12.1 (resp., $\hat{\mathcal{X}}_{v},v\in W^{\mathbf{P}}$ from Theorem 13.5) are pure with respect to the standard $t$-structure on $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{B}))$ (resp., on $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{P}))$). Evidence for their purity comes from the rank two cases, see Section 14. In Section 15, based on further evidence coming from low rank cases in ranks up to three, we state some conjectural cohomological statements that shall eventually give the description of dual exceptional collections to ${\mathcal{X}}_{v},v\in W$ in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{B})$. If true, that description must provide further evidence for purity of the objects ${\mathcal{X}}_{v},v\in W$ up to a shift.
Presumably, Theorems 11.6 and 13.3 have further extensions to larger settings than the one of semisimple algebraic groups of our paper. Much of the representation-theoretic input that we used in proving Theorems 11.6 and 13.5 also exists in the quantum setting, [7]. This suggest natural generalizations of the said theorems to the quantum case. Representation theory of parahoric subalgebras of affine Kac-Moody Lie algebras studied in [24] suggests another extension.
@@
A very interesting question is to understand a relation of our results to baric structures/staggered $t$-structures on the derived categories of equivariant coherent sheaves from [1] and [2]. Further, full exceptional collections on flag varieties appear through the computations related to the dual Steinberg basis of ${\sf K}^{0}_{\mathbf{G}}(\mathbf{G}/\mathbf{B})$, [17]. These, in turn, are related to Lusztig’s asymptotic affine Hecke algebra and to Lusztig’s canonical basis in the $\mathbf{T}$-equivariant $\sf K$-theory of Springer fibres. Ultimately, the problem is to relate the objects $X_{p}$, $p\in W$ of the present paper to Lusztig’s canonical basis.
1.4 Leitfaden
Corollary 2.21
($\mathbf{B}$-cohomology vanishing
for $P(\lambda)\otimes Q(\mu)\otimes\nabla_{\nu}$)
Theorem 4.1 (Triangularity)
Theorem 6.2
($\mathrm{R}^{i}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{G}}(P(\lambda)\otimes Q(\mu))=0$
for $i>0$)
Theorem 6.5 (6.6)
(Computing ${\mathrm{ind}}_{\mathbf{B}}^{\mathbf{G}}(P(-e_{v})\otimes Q(e_{w}))$
for $w\succeq v$)
Theorem 8.20
$\scriptstyle{Refined\ version\ of\ generation}$
Section 9.1 (“Cut at $p\in W$”)
Section 10
(Vanishing of the cone of $X_{p}\rightarrow Y_{p}$)
Section 9.2
(Construction of $X_{p}$’s and $Y_{p}$’s)
Section 11
(Constructing $\bf G$-semi-orthogonal
 decompositions of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B})$)
with the help of objects $X_{v},v\in W$
Sections 12 and 13
(Full exceptional collections
on ${\mathbb{G}/\mathbb{P}}$)
Acknowledgements
A.S. is grateful to Henning Krause for his unwavering support of this work and for his and Charles Vial’s joint invitation to the BIREP group of the University of Bielefeld for a long-term visit in 2023-2025. The research of A.S. was supported by the Deutsche Forschungsgemeinschaft (SFB-TRR 358/1 2023 - 491392403). He also acknowledges the hospitality of the Mathematical Institute of the University of Utrecht during his visits in May 2024, and in March and May 2025.
1.5 Generalities
1.5.1 Semisimple groups
@@
Let ${\mathbb{G}}\rightarrow\mathbb{Z}$ be a split simply connected semisimple Chevalley group scheme (a smooth split affine group scheme over ${\rm Spec}(\mathbb{Z})$ whose geometric fibres are connected simply connected semisimple algebraic groups), and ${\mathbb{G}}/{\mathbb{B}}\rightarrow\mathbb{Z}$ be the corresponding Chevalley flag scheme (resp., the corresponding generalized flag scheme ${\mathbb{G}}/{\mathbb{P}}\rightarrow{\mathbb{Z}}$ for a standard parabolic subgroup scheme ${\mathbb{P}}\subset\mathbb{G}$ over ${\mathbb{Z}}$). Then ${\mathbb{G}/\mathbb{P}}\rightarrow{\rm Spec}({\mathbb{Z}})$ is flat and any line bundle $\mathbb{L}$ on ${\mathbb{G}/\mathbb{P}}$ also comes from a line bundle on ${\mathbb{G}/\mathbb{B}}$.
Remark 1.6.
For the proofs we require that ${\mathbb{G}}$ is simply connected. But that is fine, because otherwise one just views the generalized flag scheme as a homogeneous space for the simply connected cover of ${\mathbb{G}}$.
@@
Given an affine algebraic group $\mathbf{H}$, flat over a noetherian base ring ${\sf k}$, an $\mathbf{H}$-module is a ${\sf k}[\mathbf{H}]$-comodule (a rational representation). The abelian category of $\mathbf{H}$-modules is denoted ${\mathrm{Rep}}(\mathbf{H})$ and ${\mathrm{rep}}(\mathbf{H})$ denotes the full abelian subcategory of representations which are finitely generated over the base ring ${\sf k}$. (cf. [5, Section 1]).
Lemma 1.7.
([32, I 2.7(4), I 4.2(1), I 4.4]) Let $M$, $N\in{\mathrm{Rep}}(\mathbf{H})$ with $M$ finitely generated and projective over ${\sf k}$. Then $M^{*}=\mathop{\mathrm{Hom}}\nolimits_{\sf k}(M,{\sf k})$ is an $\mathbf{H}$-module and
${\mathrm{Ext}}^{n}_{\mathbf{H}}(M,N)=H^{n}(\mathbf{H},M^{*}\otimes N).$
Theorem 1.8 (Universal coefficient Theorem).
Let ${\sf k}$ be a Dedekind ring and let $\mathbf{G}$ be an affine algebraic group, flat over ${\sf k}$. Let ${\sf k}^{\prime}$ be commutative ${\sf k}$-algebra and $n\geq 0$.
  1. 1.
    There is for any $\mathbf{G}$-module $M$, flat over ${\sf k}$, an exact sequence
    $0\to H^{n}(\mathbf{G},M)\otimes{\sf k}^{\prime}\to H^{n}(\mathbf{G}_{{\sf k}^{\prime}},M)\to\mathrm{Tor}_{1}^{\sf k}(H^{n+1}(G,M),{\sf k}^{\prime})\to 0$
  2. 2.
    Let $X$ be a closed reduced $\mathbb{B}_{\sf k}$-invariant subscheme, flat over ${\sf k}$, of the generalized flag variety $(\mathbb{G}/\mathbb{P})_{\sf k}$. For any $\mathbb{P}_{\sf k}$-module $N$, finitely generated and flat over ${\sf k}$, we have an exact sequence
$0\to H^{n}(X,\mathcal{L}(N))\otimes{\sf k}^{\prime}\to H^{n}(X_{{\sf k}^{\prime}},\mathcal{L}(N))\to\mathrm{Tor}_{1}^{\sf k}(H^{n+1}(X,\mathcal{L}(N),{\sf k}^{\prime})\to 0,$
where $\mathcal{L}(N)$ is the vector bundle associated to $N$ ([32, I 5.8]).
Proof.
The first part is [32, Proposition I 4.18]. The second part is a very special case, because by [32, I 5.8(2)] these sheaf cohomologies may be computed through algebraic group cohomology. The second part is tailored to our needs. One may also prove the second part with Čech cohomology. Of course we allow $\mathbb{P}=\mathbb{B}$. ∎
@@
Let $\mathbf{G}$ be obtained from ${\mathbb{G}}$ by base change along ${\rm Spec}({\sf k})\rightarrow{\rm Spec}(\mathbb{Z})$. When ${\sf k}={\mathbb{Z}}$ this just means that $\mathbf{G}=\mathbb{G}$.
@@
Now let ${\sf k}$ be a field of arbitrary characteristic. So $\mathbf{G}$ is a split semisimple simply connected affine algebraic group. Let $\mathbf{T}$ be a split maximal torus of $\mathbf{G}$, obtained by base change from a split maximal torus $\mathbb{T}$ in $\mathbb{B}$. Let $W=N_{\mathbf{G}}({\mathbf{T}})/\mathbf{T}$ be the Weyl group and let $\mathbf{B}\supseteq\mathbf{T}$ correspond with $\mathbb{B}\supseteq\mathbb{T}$. Let $X=X(\mathbf{T})$ be the weight lattice and $\Phi$ be the root system. We choose the system $\Phi^{+}$ of positive roots which makes $\mathbf{B}$ the negative Borel and let $\Pi$ be the set of simple roots. Let (,)(~{},~{})( , ) be a non-singular, symmetric, positive definite $W$-invariant form on ${\mathbb{R}}\otimes_{\mathbb{Z}}X$ and let $X(\mathbf{T})_{+}=\{\lambda\in X:(\lambda,\alpha^{\vee})\geq 0$ for all $\alpha\in\Pi$}, be the set of dominant weights, where $\alpha^{\vee}=\frac{2\alpha}{(\alpha,\alpha)}$ for $\alpha\in\Phi$.
@@
We want to emphasize once again the fact that our convention (which is that of [32]) is that ${\mathbf{B}}$ corresponds to the negative roots (as opposed to the conventions of [28], [39], [43], and [48]). For that reason, the translation of various results from loc.cit. that are used extensively in this paper requires some care.
@@
Let $\Pi=\{\alpha_{1},\dots,\alpha_{r}\}$ be the set of simple roots relative to our choice of $\Phi^{+}$. The fundamental weights $\omega_{i}$ satisfy $(\omega_{i},\alpha_{j}^{\vee})=\delta_{ij}$. A parabolic subgroup of $\mathbf{G}$ is usually denoted by $\mathbf{P}$; in particular, for a simple root $\alpha\in\Pi$, denote ${\mathbf{P}}_{\alpha}$ the minimal parabolic subgroup of $\mathbf{G}$ associated to $\alpha$.
[…]
@@
The weight lattice $X$ has a natural partial order $\geq_{d}$, known as the dominance order: for $\lambda,\mu\in X$ we write $\lambda\geq_{d}\mu$ if $\lambda-\mu$ is a sum of positive roots, with repetitions allowed [^1] That is, $\lambda-\mu$ is of the form $\sum_{\alpha\in\Phi^{+}}n_{\alpha}\alpha$, for non-negative integers $n_{\alpha}$.. We denote $w_{0}$ the longest element of the Weyl group $W$ and let $\lambda^{\ast}=-w_{0}\lambda$ for $\lambda\in X(\mathbf{T})_{+}$ (dual or contragredient, cf. [31, Section 2.2]).
@@
For $\lambda\in X$ let $\nabla_{\lambda}$ or $\nabla(\lambda)$ be the induced module ${\mathrm{ind}}_{\mathbf{B}}^{\mathbf{G}}({{\sf k}}_{\lambda})$ [32, I, Section 3.3]. It is finite-dimensional and non-zero if and only if $\lambda$ is dominant. For $\lambda\in X(\mathbf{T})_{+}$ we denote by $\Delta(\lambda)$ the Weyl module $(\nabla_{\lambda^{\ast}})^{\ast}$. Then $\nabla_{\lambda}$ has simple socle $L(\lambda)$ and $\Delta(\lambda)$ has simple head $L(\lambda)$. If ${\sf k}={\mathbb{Z}}$ then $\nabla_{\lambda}$, $\Delta(\lambda)$ are finitely generated and free over ${\sf k}$, but heads and socles make less sense.
@@
([32, II Proposition 4.13, B.4]) Let $\lambda\in X(\mathbf{T})_{+}$.
${\mathrm{Ext}}^{i}_{\mathbf{G}}({\sf k},\nabla(\lambda))={\mathrm{Ext}}^{i}_{\mathbf{B}}({\sf k},{\sf k}_{\lambda})=\left\{\begin{array}[]{l}{\sf k},\quad{\rm for}\quad i=0,\ \lambda=0,\\ 0,\quad{\rm otherwise}.\\ \end{array}\right.$
1.5.2 Triangulated categories
@@
2 Rappels: $\mathbf{B}$-modules
2.1 Highest weight category structure on ${\mathrm{Rep}}(\mathbf{G})$
@@
Recall that relative to the choice of the system of positive roots $\Phi^{+}$ there is a partial order $\leq_{d}$ on $X(\mathbf{T})$: given two weights $\lambda,\mu\in X(\mathbf{T})$ we write $\lambda\leq_{d}\mu$ if $\mu-\lambda$ is a nonnegative integer combination of elements of $\Phi^{+}$. Simple modules in ${\mathrm{Rep}}(\mathbf{G})$ are parametrized by the dominant weights $X(\mathbf{T})_{+}$, [32, II, Section 2]. By [16] and [20], there is a highest weight category structure on ${\mathrm{Rep}}(\mathbf{G})$ with the weight poset $(X(\mathbf{T})_{+},\leq_{d})$: the standard modules in this structure are the Weyl modules $\Delta_{\lambda},\lambda\in X(\mathbf{T})_{+}$ and the costandard modules are the dual Weyl modules $\nabla_{\lambda},\lambda\in X(\mathbf{T})_{+}$.